OpenStudy (anonymous):

The circle formed by the blades of a fan has a circumference of thirty-seven and five-sevenths inches. What is its diameter? Use twenty-two sevenths for pi.

OpenStudy (anonymous):

\[2\pi r=37\frac{ 5 }{ 7 }\] \[d=2r\] \[r=\frac{ 37 }{ 2(22/7) }\frac{ 5 }{ 7 }\] \[r=\frac{ 37(5) }{ 22(2) }\approx 4.2\] d=2(4.2)=8.4"

OpenStudy (anonymous):

Thank you! Thats really helpfull.

OpenStudy (anonymous):

you're welcome